OpenStudy (anonymous):

how do I solve for z? 9/25w+x=22/25z

OpenStudy (anonymous):

is this 22/25(z)

OpenStudy (anonymous):

yes a fraction

OpenStudy (anonymous):

Yes but 22 z/25 is different to 22/25 (z) so which one?

OpenStudy (anonymous):

22/25(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9/25w+x=22/25z multiply by 25z 25z (9/25w) = 22 now divide by (9/25w) 25z = 22/(9/25w) now divide by 25 z = 22/25(9/25w)

OpenStudy (anonymous):

what happen to the x?

OpenStudy (anonymous):

sorry that is my cut and paste skill for you! Wait

OpenStudy (anonymous):

9/25w+x=22/25z multiply by 25z 25z (9/25w + x) = 22 now divide by (9/25w + x) 25z = 22/(9/25w + x) now divide by 25 z = 22/25(9/25w + x)